OpenStudy (anonymous):

Simplify: (m^3n^4m^6n^2)

OpenStudy (anonymous):

use the exponential property x^a*x^b=x^(a+b) then m^3n^4m^6n^2 = m^(3+6)n^(4+2) = m^9n^6